OpenStudy (anonymous):

solve x dy/dx+3y=4x , y(0)=0 @zepdrix

OpenStudy (turingtest):

This is a linear equation and can be solved with an integrating factor. Do you know how to use that technique?

OpenStudy (anonymous):

weve gone over them in class however what we do in class seems much easier than on the homework..

OpenStudy (turingtest):

I have to eat dinner. In the meantime, use this to try to figure out the integrating factor, mu.

OpenStudy (turingtest):

oops http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (anonymous):

Hi friend, I just give the idea for you to consider, you have the equation as x y' + 3y = 4x you should change a little bit as xy' + 3y - 4x =0. it looks like the formula to figure out characteristic equation with the variable is y and its derivative. Don't take x as variable as usual. take it as a number and solve for y. since you have y(0) =0, try my way. (I'm taking discrete math, not master enough to give out the whole thing but i can see it is characteristic equation form

OpenStudy (anonymous):

took discrete. it was hard. bu this is a whole new level of math. i can see where your talking about and gave it a though however im gnna need an integrating factor, just gotta figure out which one

OpenStudy (anonymous):

thank you, I am waiting for the solution from others. I need it too. for my math classes. If you know how to solve it, post it please

OpenStudy (turingtest):

first step: get the coefficient of y' to be 1 so divide both sides by x y'+3y/x=4 the integrating factor is then\[\large\mu(x)=e^{\int\frac{dx}x}\]